# Tightrope-walking rationality in action: Feyerabendian insights for the foundations of quantum mechanics

### v2

@@ 4 Specific insights into the foundations of QM
Popper [54] had put forward his propensity-based interpretation of probabilities as a way to make them entirely objective properties of the world, thus removing subjective elements (e.g. in reference to knowledge of belief) from probabilistic theories like QM or classical statistical mechanics. It also had the goal, once this objective nature had been established and reference to ”observers” or ”agents” had been shown to be unnecessary, to support the view of QM as an incomplete description of reality, in the spirit of Einstein and against Bohr (and the whole Copenhagen ”instrumentalist” perspective). Specifically, Popper maintained the classical realist picture in terms of particle trajectories and other observables taking definite values, even if they remain ”hidden” in the ”statistical” account provided by QM. The argument by Popper (as recounted by PKF) goes roughly as follows: many apparently surprising or even paradoxical features of QM are quickly recognised, once a propensity view of probabilities is adopted, as ordinary features of probability theory; this means that QM is simply a statistical theory; as a statistical theory, QM does not account for individual events but only for ensembles, while individual processes remain ”hidden” from its description of the world, even if they remain in principle definite (as in classical mechanics); mistaking features of probability theory for features of a new mechanics of individual processes was the mistake of the founding fathers of QM, resulting from their having followed blindly a mistaken philosophy (”instrumentalism”).
@@
Upon analysis of Popper’s arguments, and of Bohr’s views, PKF concludes that the former has not offered, really, any cogent supporting argument for a classical realist picture hidden behind QM and that his criticism of Bohr’s position on the basis of his propensity view of probabilities (the starting point of his argument) simply cannot be a criticism at all.
@@
Reporting the rest of the passage already quoted above:
@@
This joint dependence of the outcome of a probability experiment on system and conditions suggests not separating the two, even in words, but regarding them as different features of a single indivisible block […] it is evident that as long as we consider probabilities a particular system cannot be separated from its surroundings, not even in thought. The probability of a die that has been shaken in a randomizer cannot therefore be explained as the result of an interaction between the die and the randomizer, or of a disturbance of the die by the randomizer. Any such account assumes that die and randomizer are separate systems which may affect each other in various ways. And while this is a perfectly adequate assumption in dynamics it must not be made in the case of probabilities which apply to the total experimental set-up. […] The reader who is familiar with complementarity will at once recognize the great similarity with the view just described. Ever since 1935 […] Bohr has emphasized the need to refer the statistical results of the quantum theory to experimental conditions and to explain some of their changes not by a causal influence, but by a change of these conditions. […] The very same dependence of statistical results on experimental conditions is emphasized, about ten years later, by Popper who writes: ’We may say that the experiment as a whole determines a certain probability distribution’ which will change or not ’relative to what we are going to regard as a repetition of our experiment; […] in other words, [probabilities] are relative to what experiments are, or are not regarded as relevant to our statistical test’ . ’The experiment as a whole’ - that is precisely what Bohr means to express by his notion of a phenomenon which is supposed to ’refer to the observations obtained under specific circumstances including an account of the whole experimental arrangement’ . […] complementarity and the propensity interpretation coincide - as far as probabilities are concerned. In both cases probabilities are properties of ’blocks’, be they now called ’experimental arrangements’, or ’phenomena’. They are not properties, not even tendencies, of individual physical systems. - [1], page 249-252
@@
Thus, PKF concludes, Popper’s starting point, albeit correct, is definitely not enough to conclude that QM is just a statistical theory (only partially accounting for an underlying classical reality) and that its apparently novel, when not paradoxical, aspects are just features of probability theory. PKF, indeed, argues correctly that, even granted the probabilistic character of QM and its relational or contextual nature (in the sense of propensities or complementarity), none of the conceptual difficulties presented to us by QM is resolved: uncertainty, superposition and quantum fluctuations, not to mention entanglement; and none of the physical arguments and experimental results that directed the Copenhagen school away a classical realist position is made less relevant or explained differently by the propensity view. Propensities can be made part of QM, and this is indeed what Bohr did, but QM still requires from us a drastic rearrangement of our picture of the world, which is what the efforts of Bohr and his colleagues struggled to achieve.
@@
It remains to say a few words about Popper’s contention that ’the reduction of the wave packet […] has nothing to do with the quantum theory: it is a trivial feature of probability theory’. The ’feature’ which Popper has in mind is of course the dependence of probabilities on experimental conditions which entails that they change abruptly when the conditions change. Now the fact that the Born Interpretation is part of the quantum theory entails that this feature is part of the quantum theory also […]. But while Popper, having before himself the ready-made result of the labours of the Copenhagen school, can regard the situation as a ’trivial feature of probability theory’, Bohr has shown how difficult it is to have both the quantum theory and a propensity interpretation of its predictions, and how many modifications are needed in order to achieve a combination of this kind […]. Now the ’reduction of the wave packet’, which leads from the unmodified quantum theory to the classical level with its propensities, contains all these modifications, including the approximations which are needed to eliminate the unwanted effects of a realistic interpretation of the principle of superposition […]. - [1], page 259-260
@@
On this point, PKF is correct, in our opinion. Funnily enough, in the same context, he is also missing another side of Popper’s position, which is however important, in modern interpretations of QM and also for appreciating better Bohr’s views. It seems to us, in fact, that in the quoted sentence, Popper had identified a key aspect of any interpretation of QM in which this is understood as being intrinsically about probabilities and in which quantum states are simply encoding of the same probabilities, devoid of an ontic status (they are not, that is, (functions of) real, intrinsic properties of the system they are associated to). In any such interpretation, the infamous ”measurement problem” is automatically solved or, better, dissolved without any need to modify or complement the QM formalism. In an epistemic interpretation of quantum states, in which they encode knowledge or information or belief about the properties of the system, the collapse of the wave function associated to measurements, thus to the acquisition of new information about the system, is simply the updating of our knowledge and beliefs. The same is true in classical statistical mechanics, where the probability distribution over the classical phase space (or configuration space) also ”collapses” to a new one upon measurement, i.e. it has to be substituted by a new one to reflect the newly acquired information. In a more objective perspective (as it is Popper’s) corresponding to a propensity interpretation of probabilities, it corresponds to an objective change of the context in which the system is defined and on which the probabilities depend, as we have seen. So, while PKF is right in stressing that the propensity view or a statistical interpretation of QM does not resolve all the conceptual puzzles it poses to our classical view of the world, and that the hard-won results of Bohr and co. need to be properly acknowledged. But he is mistakingly downplaying this key insight by Popper about a natural resolution of at least one such puzzle offered by the propensity view [^12] One could provocatively say that PKF, for once, has failed to be properly tightrope-walking between the severe critique of Popper and a proper appreciation of the insights Popper had to offer on the subject..
@@
Closer to the views of Bohr, also the so-called neo-Copenhagen perspectives on QM dissolve the measurement problem by the same interpretative step. But, just like in Bohr, as correctly noted by PKF, this step does not imply instrumentalism, but rather points to a new form of realism, radically departing from the classical one. Such neo-Copenhagen perspectives on QM include (among others) the already mentioned Relational Quantum Mechanics [40], QBism [42], the approaches of Brukner-Zeilinger [41] and Müller [39], inspired by information theory, and Healey’s pragmatism [44]. They differ in many aspects, but they can all be understood as taking onboard the key elements of Bohr’s view that had been identified by PKF, while dropping the requirement that the observation context or the observer herself have to be characterized in classical terms, as Bohr insisted. We could argue that also this last point of departure from Bohr was hinted at by PKF, when criticizing the uncritical reliance on facts of experience and on the classical language used to express them, often present in naive instrumentalism. PKF’s criticism then would extend to the requirement that the classical characterization of results of observation, and of its context should necessarily resist to a generalization in QM, or that, in other words, the new ‘quantum realism’should find an inevitable boundary in this classical characterization of the observer or the observation.
@@
More precisely, freed from this limitation, the neo-Copenhagen interpretations of QM share: a) the relative nature of quantum states, giving to them an epistemic, as opposed to ontic, characterization [^13] This could be in principle be understood in the ”hidden variables” (or ”statistical”, in the wording of Popper), as implying a non-trivial link with ontic states encoding the actual reality of physical systems. This is not the case here. No such hidden reality is assumed and while quantum states are not themselves properties of systems, they do not need to be supplemented by anything. [^14] While this seems to apply to all ‘neo-Copenhagen’ interpretations, it must be added that interpretations in this broad class differ significantly from one another. In some of them (Relational Quantum mechanics being an example) the objective, albeit relational or contextual, nature of quantum states is stressed, and the role of ‘knowledge’downplayed; the result is possibly an ontic understanding of quantum states, in which they are however not intrinsic properties of individual systems but extrinsic ones associated to pairs of systems. This is reflected also in the other aspects of the interpretations, e.g, the nature of probabilities or the nature of the ‘observing system’[55].; and thus a relative nature of quantum ‘facts’or phenomena, their context-dependence; we could label this as pointing to an epistemology of perspectival objectivity, where the context-dependence or relational aspects do not imply subjectivity but indeed a new form of (admittedly weaker) objectivity; consequently, b) a view of reality as constantly ‘moulded’by the act of observation, of interaction between observer and observed, each embodied and naturalised, but fully defined only in terms of their relations and interactions; we could label this as a metaphysics of participatory realism. Where they differ is in the articulation of what physical and conceptual features should be attributed to the ‘observer’system in the QM context, in other words the extent to which it can be embodied and naturalised and any subjective elements removed from it; related to this, they also differ in the interpretation of quantum probabilities, again ranging from a most subjective bayesian one to an objective propensity-based one. For more details about the conceptual aspects of neo-Copenhagen interpretations, and their comparison, see [55].
@@ 5 Concluding remarks
For a working theoretical physicist and a philosopher of physics, PKF is always an interesting reading, whether one agrees or not with his arguments or conclusions. The main reason is that one perceives distinctively the goal of helping scientific progress, rather that promoting one or the other philosophical perspective as such; it sets his priorities, but certainly does not imply (as it is often the case for scientists reflecting on their own theories) a lack of conceptual rigour or philosophical depth; PKF’s job, as a philosopher, was to put this rigour and depth at the service of scientific progress [^15] This is true in the context of philosophy of science, while at a more general level, scientific progress itself should be ultimately at the service of human progress, including freedom, happiness and other values, according to PKF. This a conclusion most forcefully advocated in his later production, and one basis (another being his ‘relativism’, however subtle and controversial, in epistemological matters, for limiting the power of ‘science’ and scientists in society [56].. This is also the result of his general views about the relation between physics (science, more generally) and philosophy, as we have discussed above. A key element, that we stressed several times, is the constant worry to be prey of any form of dogmatism, a defining feature of his ‘tightrope-walking rationality’and another one that is (or should be) natural to working scientists too. We hope to have made this apparent in our illustration of the many insights that PKF offered in his work on the foundations of QM, in turn mostly a case study of his general work on the realism vs instrumentalism debate, itself central in PKF’s philosophy of science. Personally, we very much value these insights.
@@
We recognize their limitations. They mostly concerned an analysis of Bohr’s views (with related discussions of the views of Bohm and other early quantum physicists) and of early puzzles of QM. They are also, to some extent, outdated. Unfortunately, PKF stopped working on the foundations of QM exactly when this area of research started becoming more fertile as a scientific domain, and we do not know what PKF thought of Bell”s theorems, or of Kochen-Specker theorem, to name some key early result in this area; and of course PKF could not comment on more recent ones, like Psi-ontology theorems, the interface between quantum foundations and quantum information, the developments of Bohmian mechanics and collapse models, or the more recent epistemic or neo-Copenhagen interpretations of QM. The same inevitable limitations concern PKF’s analysis of realism, another topic which has seen a number of new philosophical directions spurring in more recent times, including also some that could be linked to some of PKF’s insights that we discussed, e.g structural realism with its emphasis on relations, or perspectival realism. It would be unfair to focus much on these limitations. It would also be unfair, in our opinion, to emphasize the fact that PKF’s insights did not concur to build any coherent, comprehensive view on QM or realism, or that he did not complement, so to speak, his insightful criticisms with more constructive work. True, but this sounds less surprising and less disappointing in light of our previous remarks on the intended role for philosophy in its interactions with physics, according to PKF’s.
@@ Acknowledgements
We acknowledge funding from DFG research grants OR432/3-1 and OR432/4-1, as well as from FQXi grant FQXi- RFP-IPW-1908. We thank A. Barzegar, S. Hartmann and E. Curiel for useful discussions on these topics. We thank E. Margoni for a careful reading of the first version of this paper, for many useful comments, and for directing us to Hermann’s work, as well as Peter Morgan for further interesting comments. We thank the two anonymous reviewers fo their many useful comments and criticisms and for pointing out several ways in which the first version of this work could be improved. We also express our gratitude to the editors of this special issue for the invitation to contribute to it and for the remarkable patience demonstrated in dealing with us.

### v3

@@ 4 Specific insights into the foundations of QM
Popper [69] had put forward his propensity-based interpretation of probabilities as a way to make them entirely objective properties of the world, thus removing subjective elements (e.g. in reference to knowledge of belief) from probabilistic theories like QM or classical statistical mechanics. It also had the goal, once this objective nature had been established and reference to ”observers” or ”agents” had been shown to be unnecessary, to support the view of QM as an incomplete description of reality, in the spirit of Einstein and against Bohr (and the whole Copenhagen ”instrumentalist” perspective). Specifically, Popper maintained the classical realist picture in terms of particle trajectories and other observables taking definite values, even if they remain ”hidden” in the ”statistical” account provided by QM. The argument by Popper (as recounted by PKF) goes roughly as follows: many apparently surprising or even paradoxical features of QM are quickly recognised, once a propensity view of probabilities is adopted, as ordinary features of probability theory. This means that QM is simply a statistical theory. As a statistical theory, QM does not account for individual events but only for ensembles, while individual processes remain ”hidden” from its description of the world, even if they remain in principle definite (as in classical mechanics); mistaking features of probability theory for features of a new mechanics of individual processes was the mistake of the founding fathers of QM, resulting from their having followed blindly a mistaken philosophy (”instrumentalism”).
@@
Upon analysis of Popper’s arguments and Bohr’s views, PKF concludes that the former has not offered, really, any cogent supporting argument for a classical realist picture hidden behind QM and that his criticism of Bohr’s position on the basis of his propensity view of probabilities (the starting point of his argument) simply cannot be a criticism at all.
@@
From the rest of the passage already quoted above:
@@
as long as we consider probabilities a particular system cannot be separated from its surroundings, not even in thought. The probability of a die that has been shaken in a randomizer cannot therefore be explained as the result of an interaction between the die and the randomizer, or of a disturbance of the die by the randomizer. Any such account assumes that die and randomizer are separate systems which may affect each other in various ways. And while this is a perfectly adequate assumption in dynamics it must not be made in the case of probabilities which apply to the total experimental set-up. […] The reader who is familiar with complementarity will at once recognize the great similarity with the view just described. Ever since 1935 […] Bohr has emphasized the need to refer the statistical results of the quantum theory to experimental conditions and to explain some of their changes not by a causal influence, but by a change of these conditions. - [1], page 249-252
the conclusion being that Popper is repeating, in apparent opposition, what was already indicated by Bohr:
[…] The very same dependence of statistical results on experimental conditions is emphasized, about ten years later, by Popper who writes: ’[…] the experiment as a whole determines a certain probability distribution’ […] ’The experiment as a whole’ - that is precisely what Bohr means to express by his notion of a phenomenon which is supposed to ’refer to the observations obtained under specific circumstances including an account of the whole experimental arrangement’ . […] complementarity and the propensity interpretation coincide - as far as probabilities are concerned. In both cases probabilities are properties of ’blocks’, be they now called ’experimental arrangements’, or ’phenomena’. They are not properties, not even tendencies, of individual physical systems. - [1], page 249-252
@@
Thus, PKF concludes, Popper’s starting point, albeit correct, is definitely not enough to conclude that QM is just a statistical theory (only partially accounting for an underlying classical reality) and that its apparently novel, when not paradoxical, aspects are just features of probability theory. PKF, indeed, argues correctly that, even granted the probabilistic character of QM and its relational or contextual nature (in the sense of propensities or complementarity), none of the conceptual difficulties presented to us by QM are resolved: uncertainty, superposition and quantum fluctuations, not to mention entanglement; and none of the physical arguments and experimental results that directed the Copenhagen school away a classical realist position is made less relevant or explained differently by the propensity view. Propensities can be made part of QM, and this is indeed what Bohr did, but QM still requires a drastic rearrangement of our picture of the world, which is what the efforts of Bohr and his colleagues struggled to achieve.
@@
It remains to say a few words about Popper’s contention that ’the reduction of the wave packet […] has nothing to do with the quantum theory: it is a trivial feature of probability theory’. The ’feature’ which Popper has in mind is of course the dependence of probabilities on experimental conditions which entails that they change abruptly when the conditions change. Now the fact that the Born Interpretation is part of the quantum theory entails that this feature is part of the quantum theory also […]. But while Popper, having before himself the ready-made result of the labours of the Copenhagen school, can regard the situation as a ’trivial feature of probability theory’, Bohr has shown how difficult it is to have both the quantum theory and a propensity interpretation of its predictions, and how many modifications are needed to achieve a combination of this kind - [1], page 259-260
@@
On this point, PKF is correct, in our opinion. Funnily enough, in the same context, he is missing another side of Popper’s position, which is important for modern interpretations of QM and for better appreciating Bohr’s views. It seems to us, in fact, that in the quoted sentence, Popper had identified a key aspect of any interpretation of QM in which this is understood as being intrinsically about probabilities and in which quantum states are simply encoding of the same probabilities, devoid of an ontic status (they are not, that is, (functions of) real, intrinsic properties of the system they are associated to). In any such interpretation, the infamous “measurement problem” is automatically solved or, better, dissolved without any need to modify or complement the QM formalism. In an epistemic interpretation of quantum states, in which they encode knowledge or information or belief about the properties of the system, the collapse of the wave function associated to measurements, thus to the acquisition of new information about the system, is simply the updating of our knowledge and beliefs. The same is true in classical statistical mechanics, where the probability distribution over the classical phase space (or configuration space) also ”collapses” to a new one upon measurement, i.e. it has to be substituted by a new one to reflect the newly acquired information. In a more objective perspective corresponding to a propensity interpretation of probabilities (as in Popper’s), it corresponds to an objective change of the context in which the system is defined and on which the probabilities depend, as we have seen. So, while PKF is right in stressing that the propensity view or a statistical interpretation of QM does not resolve all the conceptual puzzles it poses to our classical view of the world, and that the hard-won results of Bohr and co. need to be properly acknowledged. But he is mistakingly downplaying this key insight by Popper about a natural resolution of at least one such puzzle offered by the propensity view [^11] One could provocatively say that PKF, for once, has failed to be properly tightrope-walking between the severe critique of Popper and a proper appreciation of the insights Popper had to offer on the subject..
@@
Closer to the views of Bohr, the so-called neo-Copenhagen perspectives on QM dissolve the measurement problem by the same interpretative step. But, just like in Bohr, as correctly noted by PKF, this step does not imply instrumentalism, but rather points to a new form of realism, radically departing from the classical one. Such neo-Copenhagen perspectives on QM include (among others) the already mentioned Relational Quantum Mechanics [55], QBism [57], the approaches of Brukner-Zeilinger [56] and Müller [54], inspired by information theory, and Healey’s pragmatism [59]. They differ in many aspects, but they can all be understood as taking onboard the key elements of Bohr’s view that had been identified by PKF, while dropping the requirement that the observation context or the observer herself have to be characterized in classical terms, as Bohr insisted. We could argue that also this last point of departure from Bohr was hinted at by PKF, when criticizing the uncritical reliance on facts of experience and on the classical language used to express them, often present in naive instrumentalism. PKF’s criticism would extend to the requirement that the classical characterization of results of observation, and of its context should necessarily resist to a generalization in QM, or that, in other words, the new ‘quantum realism’ should find an inevitable boundary in this classical characterization of the observer or the observation.
@@
More precisely, freed from this limitation, the neo-Copenhagen interpretations of QM share: a) the relative nature of quantum states, giving to them an epistemic, as opposed to ontic, characterization [^12] This could be in principle be understood in the ”hidden variables” (or ”statistical”, in the wording of Popper), as implying a non-trivial link with ontic states encoding the actual reality of physical systems. This is not the case here. No such hidden reality is assumed and while quantum states are not themselves properties of systems, they do not need to be supplemented by anything. [^13] While this seems to apply to all ‘neo-Copenhagen’ interpretations, it must be added that interpretations in this broad class differ significantly from one another. In some of them (Relational Quantum mechanics being an example) the objective, albeit relational or contextual, nature of quantum states is stressed, and the role of ‘knowledge’downplayed; the result is possibly an ontic understanding of quantum states, in which they are however not intrinsic properties of individual systems but extrinsic ones associated to pairs of systems. This is reflected also in the other aspects of the interpretations, e.g, the nature of probabilities or the nature of the ‘observing system’[70].; and thus a relative nature of quantum ‘facts’ or phenomena, their context-dependence; we could label this as pointing to an epistemology of perspectival objectivity, where the context-dependence or relational aspects do not imply subjectivity but indeed a new form of (admittedly weaker) objectivity; consequently, b) a view of reality as constantly ‘moulded’ by the act of observation, of interaction between observer and observed, each embodied and naturalised, but fully defined only in terms of their relations and interactions; we could label this as a metaphysics of participatory realism. Where they differ is in the articulation of what physical and conceptual features should be attributed to the ‘observer’system in the QM context. They differ in the extent to which it can be embodied and naturalised and any subjective elements removed from it. Relatedly, they differ in the interpretation of quantum probabilities, again ranging from a most subjective bayesian one to an objective propensity-based one. For more details about the conceptual aspects of neo-Copenhagen interpretations, and their comparison, see [70].
@@ 5 Concluding remarks
For a working theoretical physicist and a philosopher of physics, PKF is always an interesting read, whether one agrees or not with his arguments or conclusions. The main reason is that one perceives distinctively the goal of helping scientific progress, rather that promoting one or the other philosophical perspective as such; it sets his priorities, but certainly does not imply (as it is often the case for scientists reflecting on their own theories) a lack of conceptual rigour or philosophical depth. PKF’s job, as a philosopher, was to put this rigour and depth at the service of scientific progress [^14] This is true in the context of philosophy of science, while at a more general level, scientific progress itself should be ultimately at the service of human progress, including freedom, happiness and other values, according to PKF. This a conclusion most forcefully advocated in his later production, and one basis (another being his ‘relativism’, however subtle and controversial, in epistemological matters, for limiting the power of ‘science’ and scientists in society [71].. This is also the result of his general views about the relation between physics (science, more generally) and philosophy, as we have discussed above. A key element, that we stressed several times, is the constant worry to fall prey to any form of dogmatism, a defining feature of his ‘tightrope-walking rationality’and another one that is (or should be) natural to working scientists too. We hope to have made this apparent in our illustration of the many insights that PKF offered in his work on the foundations of QM, in turn mostly a case study of his general work on the realism vs instrumentalism debate, itself central in PKF’s philosophy of science. Personally, we very much value these insights.
@@
We recognize their limitations. They mostly concerned an analysis of Bohr’s views (with related discussions of the views of Bohm and other early quantum physicists) and of early puzzles of QM. They are also, to some extent, outdated. Unfortunately, PKF stopped working on the foundations of QM exactly when this area of research started becoming more fertile as a scientific domain, and we do not know what PKF thought of Bell’s theorems, or of Kochen-Specker theorem, to name some key early result in this area; and of course PKF could not comment on more recent ones, like Psi-ontology theorems, the interface between quantum foundations and quantum information, the developments of Bohmian mechanics and collapse models, or the more recent epistemic or neo-Copenhagen interpretations of QM. The same inevitable limitations concern PKF’s analysis of realism, another topic which has seen a number of new philosophical directions spurring in more recent times, including also some that could be linked to some of PKF’s insights that we discussed, e.g structural realism with its emphasis on relations, or perspectival realism. It would be unfair to focus much on these limitations. It would also be unfair, in our opinion, to emphasize the fact that PKF’s insights did not concur to build any coherent, comprehensive view on QM or realism, or that he did not complement, so to speak, his insightful criticisms with more constructive work. True, but this sounds less surprising and less disappointing in light of our previous remarks on the intended role for philosophy in its interactions with physics, according to PKF’s.
@@ Acknowledgements
We acknowledge funding from DFG research grants OR432/3-1 and OR432/4-1, as well as from FQXi grant FQXi- RFP-IPW-1908. We thank A. Barzegar, S. Hartmann and E. Curiel for useful discussions on these topics. We thank E. Margoni for a careful reading of the first version of this paper, for many useful comments, and for directing us to Hermann’s work, as well as Peter Morgan for further interesting comments. We thank the two anonymous reviewers and the editors for their many useful comments and criticisms and for pointing out several ways in which the first version of this work could be improved. We also express our gratitude to the editors of this special issue for the invitation to contribute to it and for the remarkable patience demonstrated in dealing with us.